# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ I Introduction
In order to match the BAO and SN data at a given redshift, we reconstruct $m_{\rm B}$ or $\mu$ of SN using a nonparametric ANN method implemented with the REFANN Python package (Wang et al., 2020). The ANN comprises a single hidden layer of 4096 neurons and is trained on the PantheonPlus and DESY5 datasets to yield a data‑driven reconstruction of $m_{\rm B}$ or $\mu$. From this reconstruction we extract $D_{\rm L}(z)$ at the redshifts of BAO measurements and compare it with the angular diameter distance $D_{\rm A}(z)$ inferred from the same BAO data to test the CDDR. The ANN reconstruction $m_{\rm B}$ for PantheonPlus and $\mu$ for DESY5 in this work are shown in Fig. 1.
This nonparametric framework does not assume any predefined functional form, providing a flexible approach for probing potential deviations from the standard cosmological model Wang et al. (2021); Liu et al. (2021); Qi et al. (2023); Liu et al. (2025b); Abedin et al. (2025b). The precision of the reconstructed $m_{\rm B}$ or $\mu$ and the width of their confidence intervals depend on hyperparameter choices, such as network architecture and training strategy, which influence the expressive power and generalization capability of ANN. Nonetheless, the ANN‑based reconstruction remains a powerful, model‑independent tool for evaluating the CDDR using BAO and SN data.
We perform Bayesian analysis using the publicly available package Cobaya [^4] https://github.com/CobayaSampler/cobaya Torrado and Lewis (2021) and assess the convergence of the Markov Chain Monte Carlo (MCMC) Lewis and Bridle (2002); Lewis (2013) chains using the Gelman-Rubin statistics quantity $R-1<0.01$ Gelman and Rubin (1992). The MCMC chains are analyzed using GetDist [^5] https://github.com/cmbant/getdist/ Lewis (2019).
III Results and discussions
@@
In this section, we present the constraints on $\eta_{0}$ and $M_{\rm B}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ parameterizations of $\eta(z)$, using the SDSS, DESI, PantheonPlus, and DESY5 data. The marginalized results with $1\sigma$ errors are summarized in Table 2. The triangular plots, along with the one-dimensional (1D) marginalized distributions of $\eta_{0}$ and $M_{\rm B}$ based on various datasets and parameterizations, are shown in Figs. 24.
In Fig. 2, we show the triangular plot of the constraint results obtained by treating $M_{\rm B}$ as a free parameter, using the SDSS, DESI, and PantheonPlus data. The constraints on $\eta_{0}$ are $0.060^{+0.120}_{-0.210}$ ($P_{1}$ model), $0.480^{+0.440}_{-0.990}$ ($P_{2}$ model), and $0.170^{+0.220}_{-0.480}$ ($P_{3}$ model), based on the DESI+PantheonPlus data. This indicates that in all parameterized models, there is no evidence for a violation of the CDDR. This result emphasizes the robustness of our CDDR analysis, and the consistency among the three parameterized models $P_{1}$, $P_{2}$, and $P_{3}$ also suggests that using multiple parameterizations of $\eta(z)$ can provide complementary insights into potential deviations from the CDDR. Similarly, the SDSS+PantheonPlus combination yields consistent results, indicating good agreement between the BAO measurements from SDSS and DESI DR2. The constraints on $M_{\rm B}$ using the DESI+PantheonPlus datasets are $-19.560^{+0.420}_{-0.290}$ ($P_{1}$ model), $-19.810^{+0.790}_{-0.570}$ ($P_{2}$ model), and $-19.650^{+0.570}_{-0.390}$ ($P_{3}$ model). The central values of $M_{\rm B}$ in all parameterized models are slightly lower than the commonly accepted value of approximately $-19.3$ mag. This slight shift may result from the inclusion of the possible deviation parameter $\eta_{0}$ in the CDDR analysis.
@@
It is worth noting that there is a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, as shown in Fig. 2. This correlation indicates that, different $M_{\rm B}$ priors will have a significant impact on our detection of potential deviations from the CDDR. Therefore, we further adopt the $M_{\rm B}^{\rm D20}$ and $M_{\rm B}^{\rm B23}$ priors to assess the potential impact of $M_{\rm B}$ on deviations from the CDDR. The 1D marginalized posterior distributions of $\eta_{0}$ for $P_{1}$, $P_{2}$, and $P_{3}$ models are shown in Fig. 3. When adopting the $M_{\rm B}^{\rm D20}$ prior with $M_{\rm B}=-19.230\pm 0.040$ mag, $P_{1}$, $P_{2}$, and $P_{3}$ models yield $\eta_{0}=-0.098\pm 0.052$, $\eta_{0}=-0.220\pm 0.113$, and $\eta_{0}=-0.162\pm 0.068$ using DESI+PantheonPlus data, respectively. All three models consistently show a deviation from CDDR at approximately $2\sigma$ confidence level, especially for the $P_{3}$ model, which has reached $2.4\sigma$. In contrast, if we adopt $M_{\rm B}^{\rm B23}$ prior with $M_{\rm B}=-19.396\pm 0.016$ mag, the constraint values of $\eta_{0}$ are $-0.035\pm 0.056$ ($P_{1}$ model), $-0.070\pm 0.124$ ($P_{2}$ model), and $-0.043\pm 0.088$ ($P_{3}$ model), suggesting that no deviation from the CDDR is found in all models. This indicates that if a fixed $M_{\rm B}$ prior is adopted, a bias of approximately 0.2 mag of $M_{\rm B}$ can lead to a deviation of $2\sigma$ in CDDR, emphasizing the necessity of analysing the impact of $M_{\rm B}$ prior on CDDR deviation. Furthermore, we can obtain that the absolute magnitude of the SN needs to be calibrated to approximately $-19.4$ mag in order for the CDDR to remain undeviation using the DESI and PantheonPlus data.
@@
Next, we present the 1D marginalized posterior distributions of $\eta_{0}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ models, obtained using the DESI+DESY5 and SDSS+DESY5 data, as shown in Fig. 4. Using the SDSS+DESY5 data, the constraint values of $\eta_{0}$ are $-0.035\pm 0.074$, $-0.080\pm 0.120$, and $-0.059\pm 0.093$ for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively. Similarly, the constraint values of $\eta_{0}$ are $-0.052\pm 0.066$ ($P_{1}$ model), $-0.090\pm 0.110$ ($P_{2}$ model), and $-0.068\pm 0.085$ ($P_{3}$ model) using DESI+DESY5 data, respectively. We find that these results consistently show no deviation from the CDDR in all parameterized models. Therefore, except for the case of adopting $M_{\rm B}^{\rm D20}$ prior, we did not detect any deviation from CDDR in other cases, indicating that the current data is well consistent with CDDR.
@@
Finally, it is necessary to compare the constraints on $\eta_{0}$ given by different observations to explore the capabilities of BAO and SN data. The constraints of $\eta_{0}$ using DESI+DESY5 are approximately 36%, 56%, and 47% higher than that obtained using the SDSS eBOSS DR16 BAO dataset combined with the Pantheon SN sample, for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively Xu et al. (2022). Similarly, the constraints are approximately 49%, 71%, 72% more restrictive than those obtained from six well-studied strong gravitational lensing systems and PantheonPlus SN data for three models Qi et al. (2025). Furthermore, this results are comparable to those inferred from compact radio quasars measurements and Pantheon SN samples Yang et al. (2024). Therefore, the latest BAO and SN data provide a more effective and accurate way to test CDDR.
@@ IV Conclusion
When $M_{\rm B}$ is treated as a free parameter, we find a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, indicating that the uncertainty in $M_{\rm B}$ directly affects the inferred CDDR deviation. Meanwhile, all three parameterizations show no evidence for a deviation from CDDR using the SDSS, DESI, and PantheonPlus data. The best-fit values of $M_{\rm B}$ are slightly lower than the commonly adopted value of $-19.3$ mag, suggesting that the inclusion of a deviation parameter in CDDR can lead to a small shift in $M_{\rm B}$ from its conventional calibration. When fixing the $M_{\rm B}^{\rm D20}$ prior ($-19.230\pm 0.040$ mag), we find a deviation from CDDR at the approximately $2\sigma$ level, especially in the $P_{3}$ model, where the deviation reaches $2.4\sigma$ level. In contrast, the $M_{\rm B}^{\rm B23}$ prior ($-19.396\pm 0.016$ mag) shows no deviation from CDDR in all models. This implies that a shift of approximately 0.2 mag in $M_{\rm B}$ can lead to a $2\sigma$ level deviation from CDDR. Finally, the joint analysis of DESI, SDSS, and DESY5 data maintains CDDR across all models.
@@ Acknowledgments
We thank Tian-Yang Sun, Sheng-Han Zhou, Jia-Le Ling, Yi-Min Zhang, and Si-Ren Xiao for their helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, 11835009, and 12205039), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the National 111 Project (Grant No. B16009).

### v2

@@ III Results and discussions
In this section, we present the constraints on $\eta_{0}$ and $M_{\rm B}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ parameterizations of $\eta(z)$, using the SDSS, DESI, PantheonPlus, and DESY5 data. The marginalized results with $1\sigma$ errors are summarized in Table 2. The triangular plots, along with the one-dimensional (1D) marginalized distributions of $\eta_{0}$ and $M_{\rm B}$, based on various datasets and parameterizations, are shown in Figs. 23 and 4.
@@
In Fig. 2, we show the triangular plot of the constraint results obtained by treating $M_{\rm B}$ as a free parameter, using SDSS, DESI, and PantheonPlus data. The constraints on $\eta_{0}$ are $0.060^{+0.120}_{-0.210}$ ($P_{1}$ model), $0.480^{+0.440}_{-0.990}$ ($P_{2}$ model), and $0.170^{+0.220}_{-0.480}$ ($P_{3}$ model), based on the DESI+PantheonPlus data. This indicates that in all parameterized models, there is no evidence for a violation of the CDDR. This result emphasizes the robustness of our CDDR analysis, and the consistency among the three parameterized models $P_{1}$, $P_{2}$, and $P_{3}$ also suggests that using multiple parameterizations of $\eta(z)$ can provide complementary insights into potential deviations from the CDDR. Similarly, the SDSS+PantheonPlus combination yields consistent results, indicating good agreement between the BAO measurements from SDSS and DESI DR2. The constraints on $M_{\rm B}$ using the DESI+PantheonPlus datasets are $19.560^{+0.420}_{-0.290}$ ($P_{1}$ model), $19.810^{+0.790}_{-0.570}$ ($P_{2}$ model), and $19.650^{+0.570}_{-0.390}$ ($P_{3}$ model). The central values of $M_{\rm B}$ in all parameterized models are slightly lower than the commonly accepted value of approximately 19.3-19.3 mag. This slight shift may result from the inclusion of the possible deviation parameter $\eta_{0}$ in the CDDR analysis.
@@
It is worth noting that there is a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, as shown in Fig. 2. This correlation indicates that different $M_{\rm B}$ priors will have a significant impact on our detection of potential deviations from the CDDR. Therefore, we further adopt the $M_{\rm B}^{\rm D20}$ and $M_{\rm B}^{\rm B23}$ priors to assess the potential impact of $M_{\rm B}$ on deviations from the CDDR. The 1D marginalized posterior distributions of $\eta_{0}$ for $P_{1}$, $P_{2}$, and $P_{3}$ models are shown in Fig. 3. When adopting the $M_{\rm B}^{\rm D20}$ prior with $M_{\rm B}=-19.230\pm 0.040$ mag, $P_{1}$, $P_{2}$, and $P_{3}$ models yield $\eta_{0}=-0.098\pm 0.052$, $\eta_{0}=-0.220\pm 0.113$, and $\eta_{0}=-0.162\pm 0.068$ using DESI+PantheonPlus data, respectively. All three models consistently show a deviation from CDDR at approximately $2\sigma$ confidence level, especially for the $P_{3}$ model, which has reached $2.4\sigma$. For further discussion on the $\sim 2\sigma$ deviation from the CDDR, please refer to Appendix B. In contrast, if we adopt an $M_{\rm B}^{\rm B23}$ prior with $M_{\rm B}=-19.396\pm 0.016$ mag, the constraint values of $\eta_{0}$ are $0.035\pm 0.056$ ($P_{1}$ model), $0.070\pm 0.124$ ($P_{2}$ model), and $0.043\pm 0.088$ ($P_{3}$ model), suggesting that no deviation from the CDDR is found in all models. This indicates that if a fixed $M_{\rm B}$ prior is adopted, a bias of approximately 0.2 mag of $M_{\rm B}$ can lead to a deviation of $2\sigma$ in CDDR, emphasizing the necessity of analysing the impact of the $M_{\rm B}$ prior on CDDR deviation. Furthermore, we can obtain that the absolute magnitude of the SN needs to be calibrated to approximately 19.4-19.4 mag in order for the CDDR to remain undeviated using the DESI and PantheonPlus data.
@@
Next, we present the 1D marginalized posterior distributions of $\eta_{0}$ for the $P_{1}$, $P_{2}$, and $P_{3}$ models, obtained using the DESI+DESY5 and SDSS+DESY5 data, as shown in Fig. 4. Using the SDSS+DESY5 data, the constraint values of $\eta_{0}$ are $035\pm 0.074, 0.0$$080\pm 0.120$, and $0.059\pm 0.093$ for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively. Similarly, the constraint values of $\eta_{0}$ are $0.052\pm 0.066$ ($P_{1}$ model), $0.090\pm 0.110$ ($P_{2}$ model), and $0.068\pm 0.085$ ($P_{3}$ model) using DESI+DESY5 data, respectively. We find that these results consistently show no deviation from the CDDR in all parameterized models. Therefore, except for the case of adopting the $M_{\rm B}^{\rm D20}$ prior, we did not detect any deviation from CDDR in other cases, indicating that the current data are well consistent with CDDR.
@@
Finally, it is necessary to compare the constraints on $\eta_{0}$ given by different observations to explore the capabilities of BAO and SN data. The constraints of $\eta_{0}$ using DESI+DESY5 are approximately 36%, 56%, and 47% higher than those obtained using the SDSS eBOSS DR16 BAO dataset combined with the Pantheon SN sample, for $P_{1}$, $P_{2}$, and $P_{3}$ models, respectively Xu:2022zlm. Similarly, the constraints are approximately 49%, 71%, and 72% more restrictive than those obtained from six well-studied strong gravitational lensing systems and PantheonPlus SN data for three models Qi:2024acx. Furthermore, these results are comparable to those inferred from compact radio quasars measurements and Pantheon SN samples Yang:2024icv. Therefore, the latest BAO and SN data provide a more effective and accurate way to test CDDR.
@@ IV Conclusion
When $M_{\rm B}$ is treated as a free parameter, we find a strong negative correlation between $\eta_{0}$ and $M_{\rm B}$, indicating that the uncertainty in $M_{\rm B}$ directly affects the inferred CDDR deviation. Meanwhile, all three parameterizations show no evidence for a deviation from CDDR using the SDSS, DESI, and PantheonPlus data. The best-fit values of $M_{\rm B}$ are slightly lower than the commonly adopted value of 19.3-19.3 mag, suggesting that the inclusion of a deviation parameter in CDDR can lead to a small shift in $M_{\rm B}$ from its conventional calibration. When fixing the $M_{\rm B}^{\rm D20}$ prior ($19.230\pm 0.040$ mag), we find a deviation from CDDR at the approximately $2\sigma$ level, especially in the $P_{3}$ model, where the deviation reaches the $2.4\sigma$ level. This deviation is due to a calibration tension between $M_{\rm B}^{\mathrm{D20}}$ and the Planck-calibrated $r_{\mathrm{d}}$, a mismatch that is closely connected to the $H_{0}$ tension. In contrast, the $M_{\rm B}^{\rm B23}$ prior ($19.396\pm 0.016$ mag) shows no deviation from CDDR in all models. This implies that a shift of approximately 0.2 mag in $M_{\rm B}$ can lead to a $2\sigma$ level deviation from CDDR. Finally, the joint analysis of DESI, SDSS, and DESY5 data maintains CDDR across all models.
@@ Acknowledgments
We thank Tian-Yang Sun, Sheng-Han Zhou, Jia-Le Ling, Yi-Min Zhang, and Si-Ren Xiao for their helpful discussions. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12533001, 12575049, 12473001, and 12205039), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the National 111 Project (Grant No. B16009).
Data Availability Statement This manuscript has no associated data. [Author’s comment: All data analysed in this work is publicly available and references have been provided in Sect. II.]
Code Availability Statement This manuscript has no associated code/software. [Author’s comment: In this work, we utilized the publicly available software packages Cobaya and GetDist, with relevant references provided in Sect. II]
Appendix A The SN Hubble diagram $D_{\mathrm{L}}(z)$ and the distance duality ratio $\eta(z)$
In this appendix, we present the results of the luminosity distance $D_{\mathrm{L}}(z)$ reconstructed from PantheonPlus (considering two different $M_{\rm B}$ priors) and DESY5 SN data, and the distance duality ratio $\eta(z)$, constructed using $D_{\rm A}$ from BAO and $D_{\rm L}$ from SN, as shown in Fig. 5. The blue data points corresponding to the combination DESI+PantheonPlus $\times$ $M_{\rm B}^{\rm D20}$ and the green data points for SDSS+PantheonPlus $\times$ $M_{\rm B}^{\rm D20}$ exhibit a noticeable deviation from the expected value $\eta(z)=1$, suggesting a potential violation of the CDDR. In contrast, the red data points corresponding to the combination DESI+PantheonPlus $\times$ $M_{\rm B}^{\rm B23}$ and the orange data points for SDSS+PantheonPlus $\times$ $M_{\rm B}^{\rm B23}$ are in better agreement with $\eta(z)=1$. These results underscore the limitations of testing the CDDR using calibration-dependent distance probes, particularly the dependence on the parameter $M_{\rm B}$. Additionally, the purple data points for DESI+DESY5 and the pink data points for SDSS+DESY5 are in close agreement, both being consistent with $\eta(z)=1$.
Appendix B A brief discussion on the $\sim 2\sigma$ deviation from the CDDR
To assess whether the $\sim 2\sigma$ departure from the CDDR when adopting the $M_{\rm B}^{\mathrm{D20}}$ prior is driven by the assumed BAO sound horizon, we repeated the analysis fixing $r_{\rm d}=138.3\ \mathrm{Mpc}$ Wang:2024rxm. The constraint results of $\eta_{0}$ are reported in Table 3. The updated fits show that, for all parametrisations and for both SDSS+PantheonPlus and DESI+PantheonPlus data, the values of $\eta_{0}$ are consistent with zero at less than the $1\sigma$ level; accordingly, the previously reported $\sim 2\sigma$ deviation disappears when a smaller $r_{\mathrm{d}}$ is adopted. This behaviour points to a calibration inconsistency between $M_{\rm B}^{\mathrm{D20}}$ and the Planck-calibrated $r_{\mathrm{d}}$, a mismatch that is also related to the $H_{0}$ tension.
Table 3: Constraint results of $\eta_{0}$ in the $P_{1}$, $P_{2}$, and $P_{3}$ models when $r_{\mathrm{d}}=138.3\ \mathrm{Mpc}$ is adopted Wang:2024rxm.
Model SDSS + PantheonPlus $\times M_{\rm B}^{\rm D20}$ DESI + PantheonPlus $\times M_{\rm B}^{\rm D20}$
$\eta_{0}$ $\eta_{0}$
$P_{1}$ model $043\pm 0.068 0.0$$037\pm 0.052$
$P_{2}$ model $080\pm 0.125 0.0$$080\pm 0.110$
$P_{3}$ model $065\pm 0.089 0.0$$068\pm 0.083$